pooja195 (pooja195):

You want to burn 380 Calories during 40 minutes of exercise.You burn about 8 calories per minute inline skating and 12 calories per minute swimming.How many minutes should you spend inline skating and swimming?

OpenStudy (anonymous):

47.5

pooja195 (pooja195):

@cvliboi that makes no sense? Im 100% thats not the right answer because theres supposed to be 2 answers

OpenStudy (anonymous):

I had x = skating and y= swimming so I took 8x + 12y =380 but where do I put the 40 minutes and how do I figure it?

OpenStudy (anonymous):

you got to find that out

pooja195 (pooja195):

Im so confused now >_<

pooja195 (pooja195):

Please dont give me the answer but guide me so that i can learn

OpenStudy (anonymous):

I had x = skating and y= swimming so I took 8x + 12y =380 but where do I put the 40 minutes and how do I figure it? NOW can you finish? (hint: you can solve the second equation for x: x = 40 - y Then substitute 40 - y for x in the first equation)

OpenStudy (unanimoose):

The set of equations should look like this. 8x + 12y = 380 x + y = 40

OpenStudy (unanimoose):

x and y are amounts of minutes and 40 is the total.

OpenStudy (unanimoose):

the combination of calories per given minutes are 8x and 12y, which total up to 380.

OpenStudy (unanimoose):

Now solve for x and y.

pooja195 (pooja195):

i got it from here thanks :)

OpenStudy (unanimoose):

Ok so: At first they say you watnt o burn 380 calories: that is your total of calories = 380 It's a 40 minute exercise: 40 minutes total = 40

OpenStudy (unanimoose):

you want to*

OpenStudy (unanimoose):

The calories you burn from skating and swimming are going towards your total calories of 380

OpenStudy (unanimoose):

So for for skating represent it at x and for swimming represent it as 12. x + y 8x + 12y 8x + 12y = 380

OpenStudy (unanimoose):

it as x*

OpenStudy (unanimoose):

x and y represent the minutes you spend skating and swimming so they are going towards total minutes: 40 x + y = 40

OpenStudy (unanimoose):

hope that helps :P

pooja195 (pooja195):

its does thanks